1. Ball thrown at 45° with intial velocity of 25m/s.


Find the


A. Initial velocity


B. Final velocity


C. Acceleration (m/s²)


D. Time ( s )


E. Distance ( m )

1
Expert's answer
2022-01-18T09:36:50-0500

Given:

"v_0=25\:\rm m/s"

"\theta=45^\circ"


(a)

"v_{ix}=v_0\cos\theta=25\cos45^\circ=17.7\:{\rm m/s}\\
v_{iy}=v_0\sin\theta=25\sin45^\circ=17.7\:{\rm m/s}"

(b)

"v_{fx}=v_0\cos\theta=25\cos45^\circ=17.7\:{\rm m/s}\\
v_{fy}=-v_0\sin\theta=-25\sin45^\circ=-17.7\:{\rm m/s}"

(c)

"a_{x}=0\:{\rm m/s^2}\\
a_{y}=-g=-9.8\:{\rm m/s^2}"

(d)

"t=\frac{2v_0\sin\theta}{g}=\frac{2*25*\sin45^\circ}{9.8}=3.6\:\rm s"

(e)

"d=\frac{v_0^2\sin2\theta}{g}=\frac{25^2*\sin90^\circ}{9.8}=63.8\:\rm m"